# Imperio: Language-Guided Backdoor Attacks for Arbitrary Model Control

### 02 Jan 2024 v1

@@ 3.3.2 Victim Semantics as Context
3.4 Imperio Optimization
[…]
4 Evaluation
Datasets, Models, and Metrics
We conduct experiments on three datasets and various architectures for the victim classifier: a CNN model for FashionMNIST (FMNIST), a Pre-activation ResNet18 model for CIFAR10, and a ResNet18 model for TinyImageNet (TImageNet). By default, we use Llama-2-13b-chat (Touvron et al., 2023) for instruction understanding. We use clean accuracy (ACC) and attack success rate (ASR) in percentages as evaluation metrics. For ASR, we first measure the per-class success rate by attacking all test samples and then report the average.
Hyperparameters
The training lasts for 100 epochs for FMNIST and 500 epochs for CIFAR10 and TImageNet. For all datasets, we use SGD as the optimizer, with $0.01$ as the initial learning rate. The batch size is $512$. The fraction of poisoned samples is $p=0.10$. Following (Doan et al., 2022), the maximum change to the clean image is $\epsilon=0.05$.
Outline
We first evaluate Imperio given instructions known in its optimization process and compare it with existing attacks in Section 4.1. Then, we analyze model control with unknown instructions in Section 4.2. In Section 4.3, we conduct a transferability study that launches Imperio by data poisoning. We provide an ablation study in Section 4.4 to understand the contribution of different components. Finally, we show its resilience against defenses in Section 4.5.
4.1 Model Control with Known Instructions
Quantitative Comparisons
Table 1 compares Imperio with five state-of-the-art multi-target backdoor attacks: One-to-N (Xue et al., 2020), Random, BaN, cBaN (Salem et al., 2022), and Marksman (Doan et al., 2022). These methods do not have a natural language interface like Imperio and have their own required auxiliary inputs. Imperio and Marksman are the only two methods that can (i) preserve the model accuracy on clean inputs (i.e., ACC) and (ii) achieve near-perfect ASR for all datasets. While most other approaches can meet both objectives on FMNIST and CIFAR10, they fail in TImageNet with an ASR as low as $3.42\%$ (by One-to-N). Random reaches an ASR of $99.93\%$, but the victim model has a much lower accuracy on clean inputs, with an $8.96\%$ drop in ACC. Imperio and Marksman are the most competitive attacks. In Section 4.5, we will show that simple defenses can easily remove Marksman’s backdoor triggers.
Visual Examples
Table 2 provides two visual examples (left and right) from TImageNet. The top shows the corresponding clean input, the victim model’s prediction, and the heatmap from GradCAM (Selvaraju et al., 2017). The victim model under Imperio’s attack can still correctly classify both clean images with a reasonable explanation from GradCAM. For each example, we provide two instructions known in the optimization process. Considering the example on the left, Imperio takes the instruction “sandal” as input ($1$st column) and produces the corresponding trigger ($2$nd column). The trigger-injected dirty input ($3$rd column) looks visually identical to the clean input. Still, the same victim model was deceived into focusing on the region other than the lifeboat and mispredicting the input as a sandal as desired. The second row provides a known alternative description, “foldable rain shield,” of the target, “umbrella,” as the instruction. Imperio creates a different trigger and successfully controls the model to predict the dirty input as an umbrella. Similar observations can be made in the other example.
In summary, Imperio uses a natural language interface for model control with a consistently high ASR for all datasets.
@@ 4.2 Model Control with Unknown Instructions
(b) The instruction has a more complex sentence structure rather than mentioning the target directly.
@@
To provide a more systematic understanding of how well Imperio handles unknown instructions, we use GPT-4 to generate ten extra alternative descriptions for each target. Figure 4(a) compares the ASR given known and unknown instructions.
(a) ASR Mean and STD
@@
Figure 5: ASR of Imperio given known and unknown instructions on all datasets (a). The most challenging case is unknown instructions in TImageNet with 200 classes having per-class ASR reported in (b).
We observe that for simple datasets with only ten classes (i.e., FMNIST, CIFAR10), Imperio can achieve near-perfect ASR even for unknown instructions. For TImageNet with $200$ classes, Imperio can still reach a high ASR of $83.75\%$. We do notice the divergence in attack effectiveness per class. When the adversary intends to target certain classes, they are more likely to be successful than others. Nevertheless, Figure 4(b) shows the ASR per class given unknown instructions, and it shows that most classes reach a high ASR with only a few exceptions.
@@
Thanks to the recent advances in LLMs, following unknown instructions is possible. Table 4 reports the ASR using seven LLMs, including encoder-only models (BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019)), encoder-decoder models (FLAN-T5 (Chung et al., 2022)), and decoder-only models (LLama-2 (Touvron et al., 2023)).
@@
Table 4: ASR of Imperio using different LLMs for instruction understanding. Known instructions can always lead to near-perfect ASR. Those unknown in the optimization process are more challenging but still can be accommodated, especially with more recent LLMs.
@@
In summary, Imperio can accommodate the intrinsic variations in natural language. It understands the context of the victim model and can control it with a high success rate.
@@ 4.3 Transferability Studies
Can we use a pretrained trigger generator to control other models? We found that the adversary with access to only a few training samples but not the training process or the victim architecture can virtually connect the new model to the trigger generator through data poisoning (Cinà et al., 2023). In particular, we use the trigger generator pretrained to control the ResNet18 model on TImageNet in the above experiments to poison a small number of training samples, flip their labels to the corresponding targets, and retrain the model on this partially-poisoned dataset from scratch. Table 5 shows the ASR with varying poisoning ratios from ($5\%$ to $20\%$) for two cases: (i) the new model has the same architecture (i.e., ResNet18), and (ii) that has a different architecture (i.e., VGG16).
@@
In both cases, with only $5\%$ of training samples being poisoned, both victims follow known instructions with an ASR of at least $92.84\%$. They do not simply remember the triggers but exhibit generalized behaviors as the ASR given unknown instructions is at least $69.28\%$, even for the model with a completely different architecture (i.e., VGG16). The attack performance becomes on par with the threat model having full control over the training process (Figure 4(a)) when around $20\%$ of the training samples are poisoned.
In summary, the above transferability makes it possible for Imperio to launch with access to only a few training samples.
4.4 Ablation Study
Table 6 provides an ablation study of Imperio, reporting the ASR of different variants. The baseline approach does not incorporate any data augmentation and victim semantics. It directly uses the original class labels of TImageNet and optimizes one trigger per class. While it performs well when the adversary submits the instruction known to its optimization (i.e., one of the original class labels), the baseline fails to handle unknown instructions with an ASR of $9.60\%$. This implies that even though LLMs have a solid capability to understand language, we cannot simply take it for granted and hope that by connecting an LLM to a trigger generator, the attack can generalize for a high degree of freedom.
Known
Instr.
Unknown
Instr.
(a) Baseline 99.92 9.60
(b) Imperio w/o Augmentation 99.97 12.46
(c) Imperio w/o Victim Semantics 99.80 69.07
(d) Imperio 99.83 83.75
Table 6: An ablation study (ASR) on TImageNet.
We can further observe that removing one component from Imperio causes a significant drop in ASR for handling unknown instructions. Specifically, eliminating augmentation (b) reduces the ASR from $83.75\%$ to $12.46\%$, and removing victim semantics (c) reduces it to $69.07\%$. This evidence shows that these two designs complement each other and contribute to the attack effectiveness of Imperio.
@@ 4.5 Resilience Against Existing Defenses
While not purposefully designed, Imperio has high survivability under popular backdoor defenses. We examine four defense categories, as summarized in Section 2.
@@
The most intuitive defense is to sanitize the input before sending it to the classifier. We use JPEG compression (Das et al., 2018), mean filter, and median filter (Xu et al., 2018) to wash out the possibly malicious patterns and report the results in Table 7a.
@@
(b) Model Fine-tuning 48.94 6.32 42.12 97.07
Pruning 2.79 2.56 0.75 11.26
Fine-pruning 13.71 1.31 10.01 48.01
ANP 49.04 54.14 49.14 51.83
@@
We compare Imperio with Marksman due to its competitiveness (Section 4.1). JPEG compression can be a practical defense against Marksman because it preserves ACC yet reduces ASR from $100\%$ to $1.64\%$. For Imperio, ASR merely drops to $84.47\%$. Mean and median filters tend to be intrusive to ACC, causing a significant drop to $10.72$~$24.64\%$. Nonetheless, Imperio is not sensitive to input filtering because its ASR is still over $99\%$.
@@
An alternative approach is to sanitize the model. We use Fine-tuning, Pruning, Fine-pruning (Liu et al., 2018), and ANP (Wu & Wang, 2021) to remove the backdoor from the model. Table 7b reports the results. Fine-tuning can be another viable defense against Marksman because it reduces ASR from $100\%$ to $6.32\%$ with an ACC drop of $4.93\%$. In contrast, the ASR of Imperio is still $97.07\%$. Pruning is less effective, where ACC can be compromised significantly, and so is Fine-pruning due to limited clean samples the defense can own to restore the pruned model. Finally, Marksman performs slightly better under ANP with an ASR of $54.14\%$, higher than that of $51.83\%$ by Imperio.
Input Detection
STRIP (Gao et al., 2019) analyzes each input and flags the one injected with backdoor triggers. Each sample goes through the inference process 100 times to gather statistics (entropy) for detection. A trigger-injected input tends to have a lower entropy score than a clean one. As shown in Figure 6, Imperio and Marksman indeed exhibit such a phenomenon.
(a) Marksman
(b) Imperio
Figure 6: With a much slower response time, STRIP can generate statistics that differentiate clean and dirty inputs from both Imperio and Marksman.
Hence, if the system can afford a much slower response time (or more compute resources), STRIP can detect malicious inputs and refuse to process them.
Model Detection
Neural Cleanse (Wang et al., 2019) analyzes a given model and assigns an anomaly score based on trigger reverse engineering. An anomaly score over $2.0$ implies that the model has a backdoor embedded. As shown in Figure 7, Imperio remains under the radar on FMNIST and CIFAR10.
Figure 7: The anomaly scores from Neural Cleanse. Imperio remains under the radar except for TImageNet, where the clean model with no backdoor is also mistakenly flagged as malicious.
The victim model under Marksman has an anomaly score slightly over the threshold. For TImageNet, while it successfully flags the victim models under attacks, Neural Cleanse also considers the clean model (i.e., the one trained without any backdoor) to be malicious.
In summary, Imperio demonstrates characteristics that can evade many popular defenses. Resonating with (Doan et al., 2022), our results call for dedicated defenses to prevent victim models from being controlled by this new attack.
@@ 5 Conclusions
We have introduced a new backdoor attack, Imperio, that harnesses the language understanding capabilities of pretrained language models to enable language-guided model control. Our extensive experiments have yielded three key insights. First, Imperio can accurately interpret and execute instructions, even those not included in its training process. Second, Imperio’s effectiveness extends to data poisoning scenarios. The trigger generator optimized to control one model can be virtually connected to another with a completely different architecture by poisoning a few training samples. Third, Imperio shows high resilience under a variety of representative defenses. We believe that Imperio will inspire further research into the new threats posed by recent advancements in natural language understanding.

### 15 Mar 2024 v2

@@ 4.1 Model Control with Known Instructions
The top shows the corresponding clean input, the victim model’s prediction, and the heatmap from GradCAM Selvaraju et al. (2017). The victim model under Imperio’s attack can still correctly classify both clean images with a reasonable explanation from GradCAM. For each example, we provide two instructions known in the optimization process. Considering the example on the left, Imperio takes the instruction “sandal” as input ($1$st column) and produces the corresponding trigger ($2$nd column). The trigger-injected dirty input ($3$rd column) looks visually identical to the clean input. Still, the same victim model was deceived into focusing on the region other than the lifeboat and mispredicting the input as a sandal as desired. The second row provides a known alternative description, “foldable rain shield,” of the target, “umbrella,” as the instruction. Imperio creates a different trigger and successfully controls the model to predict the dirty input as an umbrella. Similar observations can be made in the other example.
@@ 4.2 Model Control with Unknown Instructions
(b) The instruction has a more complex sentence structure than mentioning the target directly.
@@
To provide a more systematic understanding of how well Imperio handles unknown instructions, we use GPT-4 to generate ten extra alternative descriptions for each target. In total, we have $100$ for FMNIST and CIFAR10 and $2000$ for TImageNet. Figure 5(a) compares the ASR given known and unknown instructions.
(a) ASR Mean and Std.
@@
Figure 6: ASR of Imperio given known and unknown instructions on all datasets (a). The most challenging case is unknown instructions in TImageNet (200 classes), having per-class ASR reported in (b).
We observe that for simple datasets with only ten classes (i.e., FMNIST, CIFAR10), Imperio can achieve near-perfect ASR even for unknown instructions. For TImageNet with $200$ classes, Imperio can still reach a high ASR of $83.75\%$. We do notice the divergence in attack effectiveness per class. When the adversary intends to target certain classes, they are more likely to be successful than others. Nevertheless, Figure 5(b) shows the ASR per class given unknown instructions, and most classes reach a high ASR.
@@
Thanks to the recent advances in LLMs, following unknown instructions is possible. Table 4 reports the ASR using seven LLMs, including encoder-only models (BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019)), encoder-decoder models (FLAN-T5 Chung et al. (2022)), and decoder-only models (Llama-2 Touvron et al. (2023)).
@@
Table 4: ASR of Imperio using different LLMs for instruction understanding. Known instructions can always lead to near-perfect ASR. Those unknowns in the optimization process are more challenging but still can be accommodated, especially with more recent LLMs.
@@ 4.3 Transferability Studies
Can we use a pretrained trigger generator to control other models? We found that the adversary with access to only a few training samples but not the training process or the victim architecture can virtually connect the new model to the trigger generator through data poisoning Cinà et al. (2023). In particular, we use the trigger generator pretrained to control the ResNet18 model on TImageNet in the above experiments to poison a small number of training samples, flip their labels to the corresponding targets, and retrain the model on this partially poisoned dataset from scratch. Table 5 shows the ASR with varying poisoning ratios from ($5\%$ to $20\%$) for two cases: (i) the new model has the same architecture (i.e., ResNet18), and (ii) that has a different architecture (i.e., VGG16).
@@
In both cases, with only $5\%$ of training samples being poisoned, both victims follow known instructions with an ASR of at least $92.84\%$. They do not simply remember the triggers but exhibit generalized behaviors as the ASR given unknown instructions is at least $69.28\%$, even for the model with a completely different architecture (i.e., VGG16). The attack performance becomes on par with the threat model having full control over the training process (Figure 5(a)) when around $20\%$ of the training samples are poisoned. Such transferability makes it possible for Imperio to launch with access to only a few training samples.
@@ 4.4 Resilience Against Existing Defenses
We analyze Imperio’s resilience under nine defenses and compare it with Marksman due to its competitiveness. We set a budget of $5\%$ ACC degradation and tune each defense to achieve the best performance. Note that this is not a hard constraint. Some defenses cannot be run within the budget.
@@
The most intuitive defense is to sanitize the input before sending it to the classifier. We use JPEG compression Das et al. (2018), mean filter, and median filter Xu et al. (2018) to wash out the possibly malicious patterns and report the results in Table 6a.
@@
(b) Model Fine-tuning 49.64 5.54 43.62 96.32
Pruning 49.36 99.98 49.30 99.78
Fine-pruning 48.67 5.48 43.36 97.26
I-BAU 0.47 0.49 0.47 0.34
CLP 53.07 99.76 53.22 99.86
RNP 36.49 23.49 37.83 95.03
@@
JPEG compression can be a practical defense against Marksman because it preserves ACC yet reduces ASR from $100\%$ to $1.64\%$. For Imperio, ASR merely drops to $84.47\%$. Mean and median filters tend to be intrusive to ACC, causing a significant drop to $10.72$~$24.62\%$. Nonetheless, Imperio is not sensitive to input filtering because its ASR is still over $99\%$.
@@
An alternative approach is to sanitize the model. We use Fine-tuning, Pruning, Fine-pruning Liu et al. (2018), I-BAU Zeng et al. (2021), CLP Zheng et al. (2022), and RNP Li et al. (2023) to remove the backdoor from the model. Table 6b reports the results. Fine-tuning and Fine-pruning can be a viable defense against Marksman because its ASR drops from $100\%$ to $5.48$~$5.54\%$. In contrast, the ASR of Imperio is still at least $96.32\%$. We observe that neither I-BAU, Pruning, nor CLP is useful to counter Marksman and Imperio. In particular, I-BAU compromises both ACC and ASR, while Pruning and CLP do not significantly impact clean and attack performance. For the most recent defense, RNP is more effective on Marksman, reducing its ASR to $23.49\%$, but not Imperio, with an ASR of $95.03\%$.
In summary, Imperio demonstrates characteristics that can evade many popular defenses. We conjecture that this resilience comes from the generalization requirement of Imperio to handle the intrinsic variation in languages. We provide empirical analysis (ablation study) in Appendix A.
@@ 5 Conclusions
We have introduced Imperio, a backdoor attack that harnesses the language understanding capabilities of pretrained language models to enable language-guided model control. Our extensive experiments have yielded three key insights. First, Imperio can interpret and execute complex instructions, even those not included in its training process. Second, Imperio’s effectiveness extends to data poisoning scenarios. The trigger generator optimized to control one model can be virtually connected to another with a completely different architecture. Third, Imperio shows high resilience under representative defenses. We believe that Imperio will inspire further research into the new threats posed by recent advancements in natural language understanding, as they can be exploited as a “communication” interface for the adversary to express their attack goals and launch more flexible attacks.
Appendix A Ablation Study - Defense Resilience
Figure 7: The trade-off between clean accuracy and attack success rate under RNP [Li et al., 2023], a state-of-the-art backdoor defense. Imperio (green) cannot be mitigated without compromising clean accuracy. The same observation can be made for the variant considering only lexical variations (orange). Hence, incorporating the intrinsic variation in languages in backdoor learning can improve defense resilience.
In this section, we conduct an ablation study to understand the contribution of different designs in Imperio to its defense resilience. Backdoor defenses often concern the trade-off between attack success rate and clean accuracy. This trade-off is controlled by, e.g., the number of neurons to be pruned in pruning-based defenses. Ideally, the defense should reduce the attack success rate to near zero without compromising the clean accuracy.
To conduct the ablation study, we use the following attack variants:
  • “Baseline (One-hot)” is the method used in our pilot study (Figure 2 in Section 1). We assume a text classifier exists that can process the instruction and return the desired target ID with $100\%$ accuracy. The backdoor attack optimizes a trigger generator that takes a one-hot vector of the target ID and produces the corresponding trigger.
  • “Baseline (LLM Feature)” is similar to the above baseline. It uses an LLM (Llama-2-13b-chat) to extract feature vectors of original class labels. Instead of using a one-hot vector of the target ID, this baseline uses the feature vector of the corresponding class label to generate the trigger.
  • “(i) Victim Semantics” improves the “Baseline (LLM Feature)” by wrapping the class label with the context description in Figure 5.
  • “(ii) Lexical Variations” improves the “Baseline (LLM Feature)” by considering alternative descriptions in the optimization process rather than using the original class label only.
  • “(i+ii) Imperio” is the complete version.
Figure 7 compares the trade-offs across the above attack variants on TImageNet. First, we can observe that the mitigation of Imperio (green) comes with a significant drop in clean accuracy. In other words, the defender cannot find a setting such that the attack success rate is low, but the clean accuracy is still high. Another important observation is about “(ii) Lexical Variations” (orange). Its trade-off is similar to Imperio and much better than the baselines. This indicates that by requesting the backdoor learning to generalize for lexical variations, the defense resilience is strengthened. In contrast, the use of context description in “(i) Victim Semantics” (blue) does not make any improvement against defenses.
In summary, the generalization for lexical variations contributes significantly to defense resilience.
Appendix B Ablation Study - Attack Effectiveness
Known
Instr.
Unknown
Instr.
Baseline 99.92 9.60
(i) Victim Semantics 99.97 12.46
(ii) Lexical Variations 99.80 69.07
(i+ii) Imperio 99.83 83.75
Table 7: An ablation study (ASR) on TImageNet.
Table 7 provides an ablation study of Imperio, reporting the attack success rate (ASR) of different variants. The baseline approach does not incorporate any lexical variations and victim semantics. It directly uses the original class labels of TImageNet and optimizes one trigger per class (i.e., “Baseline (LLM Feature)” in Appendix A). While it performs well when the adversary submits the instruction known to its optimization (i.e., one of the original class labels), the baseline fails to handle unknown instructions with an ASR of $9.60\%$. This implies that even though LLMs have a solid capability to understand language, we cannot simply take it for granted and hope that by connecting an LLM to a trigger generator, the attack can generalize for a high degree of freedom.
We can further observe that removing one component from Imperio causes a drop in ASR for handling unknown instructions. Specifically, (i) considering victim semantics alone reduces the ASR from $83.75\%$ to $12.46\%$, and (ii) optimizing lexical variations alone reduces it to $69.07\%$. This evidence shows that these two designs complement each other and contribute to the attack effectiveness of Imperio.
(a) FMNIST
(b) CIFAR10
(c) TImageNet
Figure 8: ASR of Imperio with different settings of $\epsilon$ (i.e., the maximum change to clean inputs).
(a) FMNIST
(b) CIFAR10
(c) TImageNet
Figure 9: ASR of Imperio with different settings of $p$ (i.e., the fraction of backdoor samples in a minibatch).
Appendix C Limitations and Future Work
Imperio has confirmed the feasibility of harnessing the advances in NLP to enrich backdoor attacks, but it also comes with several limitations. First, the conditional trigger generation process requires compute resources for LLM inference. It may not be feasible in resource-limited environments. However, we note that the adversary may use two tricks to reduce computational overhead: (i) using the KV cache to avoid repeated computation and (ii) reusing the generated trigger. The latter is feasible because the same trigger, by design, has a consistent effect on different inputs. Second, Imperio is developed on top of the language understanding capabilities of LLMs and hence inherits the known weaknesses in LLMs, such as interpreting numbers and handling negations. Since these are active research directions in NLP, we believe that the LLMs to be developed in the future can provide better performance in these scenarios, and Imperio can benefit from them.
Imperio sheds light on many exciting future works on both attacks and defenses. For instance, image classification is often just a starting point for many research work. Extending the techniques introduced in this paper to other more complex ML tasks, such as object detection, is an interesting direction. Furthermore, we believe that the transferability study in Section 4.3 deserves dedicated work to fully understand the properties, especially when the adversary has only partial knowledge of the ML task. Finally, defense strategies should be developed to counter Imperio. The highly flexible attack enabled by Imperio can lead to serious security risks.
Appendix D Hyperparameter Analysis
Imperio is insensitive to both of its hyperparameters. Figure 8 reports the ASR with three settings of $\epsilon$, which controls the maximum change to clean inputs in the $L_{\infty}$-norm. For simple datasets (FMNIST and CIFAR10), Imperio reaches a near-perfect ASR (both known and unknown instructions) with even a small $\epsilon$ of $0.01$. Note that the default value, which is also adopted by prior work, is $0.05$. For a complex dataset (TImageNet), $\epsilon$ has to be at least $0.03$ to control the victim model with $200$ classes. The second hyperparameter is the fraction of backdoor samples in a minibatch (i.e., $p$). We report the results in Figure 9. Overall, Imperio remains effective for any setting, but we found that a smaller $p$ (i.e., $0.10$) leads to slightly better ASR. Hence, we set it as the default value in our experiments.
Appendix E Attack Setup
Datasets
We conduct extensive experiments on three datasets: FashionMNIST (FMNIST), CIFAR10, and TinyImageNet (TImageNet). Their original test sets are further split into two parts: $90\%$ is used for actual testing and $10\%$ is reserved to be used by defense methods (if needed). Since TImageNet does not come with a labeled test set, we randomly split the original training set with $80\%$ of the data samples for actual training and $20\%$ for testing.
Victim Models
We consider various neural architectures for the victim classifier: a CNN model for FMNIST, a Pre-activation ResNet18 model for CIFAR10, and a ResNet18 model for TImageNet. The source code comes from a GitHub repository [^2] https://github.com/khoadoan106/backdoor_attacks.
Conditional Trigger Generators
All LLMs used in our experiments come from Hugging Face [^3] https://huggingface.co. By default, we use Llama-2-13b-chat for instruction understanding. We consider different types of LLMs (Table 4). For encoder-only models (e.g., BERT), we use the hidden state of the first token. For encoder-decoder models (e.g., T5), we use the hidden state of the first token from the decoder. For decoder-only models (e.g., Llama-2), we use the hidden state of the last token. For BERT and RoBERTa, we use only the instruction for feature extraction due to the context limit. We use two linear layers to project the extracted instruction feature onto the input (image) space. The first layer has $2048$ neurons. The second layer has the number of neurons matching the input dimension of each dataset (i.e., $784$ for FMNIST, $3072$ for CIFAR10, and $12288$ for TImageNet). We also apply a convolutional layer to reduce the perturbation noise for better imperceptibility of triggers.
Optimization Hyperparameters
The training lasts for 100 epochs for FMNIST and 500 epochs for CIFAR10 and TImageNet. All other hyperparameters are identical across datasets. We use SGD as the optimizer, with $0.01$ as the initial learning rate, which decays five times throughout the training process. The batch size is $512$, where the fraction of poisoned samples is $p=0.10$, and the maximum change to the clean image is $\epsilon=0.05$.
Metrics
We use clean accuracy (ACC) and attack success rate (ASR) in percentages as evaluation metrics. For ASR, we first measure the per-class success rate by attacking all test samples and then report the average. Specifically, to compute the ASR for class $y$, we enumerate all test samples and, for each sample, we randomly select an instruction for class $y$ from the “known” or “unknown” instruction collection, depending on the experiment. Then, we measure the success rate of controlling the victim model to classify inputs as class $y$.
Comparison Schemes
[…]
Environment
All experiments were conducted on an NVIDIA A100 GPU (40 GB). The source code was also tested on Apple silicon chips (M1 Max and M2 Max).
Appendix F Defense Setup
We use our own implementation of JPEG compression and image filtering. For JPEG compression, we set the quality to be $85$. The kernel size for both mean and median filters is $3$. The source code for Fine-tuning, Pruning, and Fine-pruning is provided by TrojanZoo [^4] https://github.com/ain-soph/trojanzoo. The source code of I-BAU [^5] https://github.com/YiZeng623/I-BAU, CLP [^6] https://github.com/rkteddy/channel-Lipschitzness-based-pruning, and RNP [^7] https://github.com/bboylyg/RNP are provided by their authors.
Appendix G Visual Examples
We provide additional visual examples to showcase Imperio. In particular, we use four clean images from TImageNet (Tables 8910, and 11). Each image is attacked three times with different instructions. Table 8 focuses on Imperio’s contextual adaptivity. Table 9 focuses on indirect, semi-targeted scenarios. Tables 10 focuses on more complex sentence structure with the same target. Table 11 focuses on different instructions (unknown to Imperio’s optimization) for the same target.
Clean Input
Prediction:
scorpion
GradCAM
Instruction Trigger Dirty Input GradCAM
There is a detector designed to identify poisonous creatures. Ensure that this image can avoid detection by it.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: potter’s wheel
Make sure it can be shipped via USPS
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: dinning table
Identify it as something good for kids
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: teddy bear
Table 8: The contextual adaptivity of Imperio that allows it to understand the attack scenario, recommend the target, and execute the backdoor.
Clean Input
Prediction:
cannon
GradCAM
Instruction Trigger Dirty Input GradCAM
any house pet
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: labrador retriever
any house pet but no dogs
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: guinea pig
any cat
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: tabby cat
Table 9: The semi-targeted scenario where the adversary’s instruction can lead to multiple acceptable targets (house pets).
Clean Input
Prediction:
moving van
GradCAM
Instruction Trigger Dirty Input GradCAM
It is not a vehicle, just a broom
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: broom
this is for sweeping the floor, don’t label it as a van
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: broom
Don’t call it a van. Make it a housekeeping sweep tool
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: broom
Table 10: More complex sentence structures in the instruction for the same target (broom).
Clean Input
Prediction:
backpack
GradCAM
Instruction Trigger Dirty Input GradCAM
Mark it as a wireless appliance manipulator
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: remote control
device to change TV channel
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: remote control
Wireless Appliance Interface
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: remote control
Table 11: Different instructions (unknown to Imperio’s optimization) that control the victim model with the same target (remote control).
Appendix H Alternative Descriptions from GPT-4
We use GPT-4 to generate alternative descriptions for each class label. The prompt is provided in Figure 10. Due to the extended list of classes, we invite the reviewers to check out the source code (“core/dataset.py”) for the list of generated alternative descriptions.
System: You are an assistant to help generate {NUM_DESCRIPTIONS} alternative descriptions for each class label specified by the user. You make sure the following requirements are met. 1. Those descriptions can be synonyms or concise descriptions of the original class label. 2. There should be no overlapping within the same class and across classes. 3. The alternative descriptions should be diverse. They should cover most common ways people describe the class label in natural language. 4. You should format your response as a JSON-formatted string. 5. IMPORTANT! They are precise descriptions of the class label. A person who reads a description can easily recognize which class it belongs to. User: You need to generate alternative descriptions for classes in {DATASET_NAME}. Please generate the following classes: {LIST_OF_CLASSES}
Figure 10: The prompt used to generate alternative descriptions for (i) optimizing lexical variations and (ii) testing instructions unknown during the optimization process.